OpenStudy (anonymous):

Which of the following is the solution for the inequality below>>>>-3x+2<8 A)x>-3 B)x>-2 C)x<-2 D)x<-3

OpenStudy (anonymous):

c

OpenStudy (anonymous):

its b :)

OpenStudy (anonymous):

Ook then..

OpenStudy (anonymous):

I have no idea now lol

OpenStudy (anonymous):

its b :)

OpenStudy (anonymous):

-3(-2)=8 Therefore, x must be less than -2

OpenStudy (kendricklamar2014):

Please don't give Direct Answers @Adam79 and @lizzybeth33

OpenStudy (anonymous):

Okay

OpenStudy (kendricklamar2014):

@M_palomares Let's solve this together: First we subtract both sides by 2\[-3x + 2 - 2 < 8-2\] Which equals:\[-3x < 6\]

OpenStudy (kendricklamar2014):

Second, we divide both sides by -3:\[\left(\begin{matrix}-3x \\ -3\end{matrix}\right) < \left(\begin{matrix}6 \\ -3\end{matrix}\right)\]

OpenStudy (kendricklamar2014):

Which equals:\[x > - 2 ~ (Choice B)\]

OpenStudy (anonymous):

Ok tysm:)

OpenStudy (kendricklamar2014):

No Problem :)